OpenStudy (anonymous):

morning to all doing factor polynomials and needs some help. I am trying to get this but it is not being easy. It is asking to find the common factor of 2x^2 +4x I think it is 2,x,2x,x^2 but I am not sure

OpenStudy (anonymous):

There is a 2 and and x in both terms. Observe: 2x^2 + 4x = (2)(x)(x) + (2)(2)(x)

OpenStudy (anonymous):

here are ,my choices1,2,2x,2x^2 or 2,x,2x,x^2 or 2,2x,4x,x^2

OpenStudy (anonymous):

2x^2 + 4x = (2)(x)(x) + (2)(2)(x) = (2x)(x) + (2x)(2) = (2x)(x + 2). Fully factored now.